OpenStudy (anonymous):

Express the complex number in trigonometric form. 5 - 5i

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

a+bi <---> r (cos theta+i sin theta) \(r=\sqrt{a^2+b^2} \\ \theta= \tan^{-1}(b/a)\) here, a = +5, b=-5

OpenStudy (anonymous):

r = sqrt(25 + 25), so r = sqrt 50 θ = tan^-1(-5/5) what else do i do, hart?

hartnn (hartnn):

simplify.... sqrt 50 =... ? tan^-1 (-1) = ...?

OpenStudy (anonymous):

r = 5sqrt2 θ = -pi/4

hartnn (hartnn):

correct! :) lastly, cos (-pi/4) = cos (pi/4) and sin (-pi/4) = - sin (pi/4) so, 5 sqrt 2 (cos pi/4 - i sin pi/4) got this ?

OpenStudy (anonymous):

5sqrt2(cos(pi/4) - i*sin(pi/4) I got it, but the options are 5sqrt2(cos(5pi/4) - i*sin(5pi/4) or 5sqrt2(cos(7pi/4) - i*sin(7pi/4)

hartnn (hartnn):

ok, -pi/4 = -pi/4 +2pi = 7pi/4 so, 5 sqrt 2 (cos 7pi/4 + i sin 7pi/4)

OpenStudy (anonymous):

Wow, thanks hart. you're a beast

hartnn (hartnn):

not actually, but thanks for the compliment :) welcome ^_^